# Language Models Can Predict Their Own Behavior

### v1

@@ Abstract
Autoregressive Language Models output text by sequentially predicting the next token to generate, with modern methods like Chain-of-Thought (CoT) prompting achieving state-of-the-art reasoning capabilities by scaling the number of generated tokens. However, are there times when we can infer how the model will behave (e.g. abstain from answering a question) early in the computation, making generation unnecessary? We show that internal representation of input tokens alone can often precisely predict, not just the next token, but eventual behavior over the entire output sequence. We leverage this capacity and learn probes on internal states to create early warning (and exit) systems. Specifically, if the probes can confidently estimate the way the LM is going to behave, then the system will avoid generating tokens altogether and return the estimated behavior instead. On 27 text classification datasets spanning five different tasks, we apply this method to estimate the eventual answer of an LM under CoT prompting, reducing inference costs by 65% (average) while suffering an accuracy loss of no more than 1.4% (worst case). We demonstrate the potential of this method to pre-emptively identify when a model will abstain from answering a question, fail to follow output format specifications, or give a low-confidence response. We explore the limits of this capability, showing that probes generalize to unseen datasets, but perform worse when LM outputs are longer and struggle to predict properties that require access to knowledge that the models themselves lack. Encouragingly, performance scales with model size, suggesting applicability to the largest of models.  [^1] https://github.com/DhananjayAshok/LMBehaviorEstimation
@@ 1 Introduction
Language Models trained to predict the next token of a sentence on large-scale data [58, 14] are unrivalled in general Natural Language Understanding (NLU). During inference, these models generate remarkably high-quality outputs by sequentially predicting the most appropriate output tokens to generate. Frontier LM systems of today build on this by scaling inference i.e. increasing the number of tokens or sequences generated [64, 67, 48, 20], to provide even better responses.
In this paper we ask whether LMs have learned representations that contain information, not just about the next token the model will predict, but the entire sequence as a whole. We conclude with an affirmative answer, showing that the internal states of the LM can often be used to preemptively identify how an LM will behave. This information becomes accessible before the LM has generated a single token, potentially obviating the need for further inference on those instances.
@@
Concretely (Figure 1), we train linear classifiers (probes) [3] that use the internal representation of the input tokens to predict the eventual behavior of the LM (e.g. whether an LM will comply with a request or abstain). We calibrate the probes using methods from conformal prediction [55], allowing them to estimate behavior only when confident. During inference, probes predict only when there is a provable guarantee on the estimation error, ensuring precise early warning signals for various model behaviors.
@@
We apply this method to an LM that tackles text classification by generating several explanation tokens before a class prediction (CoT). The probes estimate the final CoT prediction and exit early if confident in their estimate. On 27 datasets, spanning the tasks of MCQA, Sentiment Analysis, Topic Classification, Toxicity Detection and Fact Verification, our method reduces inference costs by 65% on average, while suffering accuracy losses of no more than 1.4%. Furthermore, on 14/27 of the datasets, our method reduces inference costs by 63% on average with no cost to accuracy. The probes generalize to unseen datasets, reducing the inference cost of CoT on OpenBookQA [44] by 68% with minimal loss to accuracy (0.4%), despite only training on other MCQA datasets.
Beyond accelerating text classification, probes can pre-emptive identify several behaviors of interest. We apply our method to a QA system that verbalizes some reasoning about the input question and then either answers or abstains if the question is deemed to be unanswerable. The probes, which are trained to estimate whether the LM will abstain, make confident predictions on over 15% of the test instances and achieve an estimation consistency of over 90% when doing so.
Probes can provide precise early warning signals for degenerate behavior (whether the model will fail to follow a specified output format) and even ‘self-reflective" properties which cannot be inferred from the text output alone (whether the model will output a high perplexity answer to a question). While the number of instances where the probes give confident predictions varies significantly across datasets and tasks, the probe maintains high consistency on confident predictions.
@@
We show that often fewer than a thousand training instances are sufficient for the probes to attain high estimation consistency, while ablations on the probing layer present a more complicated, task-specific story. Encouragingly, increasing the scale of the LM improves the performance of the probe, suggesting the method may scale favorably with ever-increasing model sizes [10, 24, 11, 16].
The probes are limited in the behaviors they can preemptively detect. While probes can identify various model behaviors, they fail on those that cannot be identified without knowledge external to the model (e.g. identifying whether the model is likely to output an incorrect answer to an MCQ instance). Additionally, probes are less confident on instances which trigger longer outputs.
@@
Figure 1: Overview of our method: During training we collect the hidden states of the input token and eventual output behavior on a variety of prompts, using them to train a probe on the hidden states. With a validation set, we calibrate the probes for conformal prediction by identifying a threshold for confident predictions. During inference, if the probes are confident in their estimate of the eventual LM behavior, we stop all computation and return the prediction.
@@ 2 Background
Hidden State Probing: Lightweight probes have long been used to interpret the internal activations of neural networks [3] and language models [50, 7]. Given a set of input prompts, we compute the generated outputs and store the internal activations computed during the forward passes of the LM (e.g. the outputs of a specific Transformer [62] layer). The instances are then assigned a classification label meant to capture some property of the input and output. The probes are then trained to predict the label using only the internal activations as input. For example, Azaria and Mitchell [7] collect the activations of the middle layer of an LM when processing the output tokens, and manually label the generations as true or false. They then train a small neural network which can identify whether the model output is truthful given the internal activations of the generated tokens. Our work focuses on behaviors that can be trivially inferred from the output text alone (e.g. infer the answer from the output, infer if the output follows a specified format) and hence does not require manual labelling. Standard practice in probing uses the internal states of all tokens, even generated ones. We take an alternate approach, training probes on the internal states of the input tokens alone, allowing them to predict properties of the output tokens before they have been generated.
@@
At inference time, the prediction set for a test instance $\mathbf{x}$ is simply $\{\hat{c}|S(\mathbf{x},\hat{c})\leq q_{\alpha}\}$.
For the score function provided above, no two classes can achieve a value greater than 0.5 simultaneously, implying that with $\alpha>0.5$ the prediction set at inference time is either empty (deferral/abstention) or consists of a single class label.
@@ 3 Can Internal States Predict Eventual Behavior?
Figure 2: Left: Estimation consistency (accuracy of predictions on the answer that CoT will eventually select. Right: Conformal estimation of CoT answer. Probes are often confident on sizeable portions of the test sets (37% on average) with low error even on datasets with poor total consistency.
In this section, we ask whether the internal states of Language Models contain information on the properties of the output tokens the model will eventually generate. We explain our methodology with the running example of a Chain-of-Thought prompted model answering a multiple choice question.
The CoT model is expected to reason about the input question before giving its final answer, a paradigm that has been shown to enable superior performance on a wide range of reasoning tasks [64, 36, 63]. Suppose we are given an input prompt of a question with two answer options:
   Question: Osteoclasts
   Option A: deposit bone and differentiate from monocytes.
   Option B: resorb bone and differentiate from monocytes.
   Explanation:
Llama3.1-8B [16] responds to this with
Osteoclasts are cells that break down bone tissue and are responsible for
bone reabsorption. They differentiate from monocytes and macrophages.
Answer: B
We can reasonably infer that the model will select option B by the time it generates the token ‘reabsorption’, but the tokens before that are generic enough that they do not eliminate either option. However, is it possible that the model contains signals that can identify the answer by the time it generates its initial output token ‘Osteoclasts’, and simply takes multiple tokens to express it in an interpretable way?
We collect the output of the 18th (18/32) Transformer layer of Llama3.1-8B when processing the final token of the input (i.e. the final ‘:’), and train linear classifiers to predict what the eventual answer prediction will be. We ablate layers in Section 6 and reproduce results on other LMs in Appendix A.
The probes achieve (Figure 2) higher than random consistency (accuracy on prediction of model behavior) on all datasets, suggesting that input embeddings contain information on the eventual answer the LM will output. However, the performance varies significantly across datasets, suggesting that a naive attempt to exit early with these estimations could result in catastrophic failure.
@@ 3.1 Creating robust behavior estimators with conformal prediction
Unless one believes that the explanation generated by the LM has no causal influence on the eventual answer, there will always be instances where there is not enough information in the internal state of the input tokens to conclude what the future model behavior will be. The ideal behavior estimation system handles such cases, making consistent predictions when it is confident and deferring otherwise.
Hoping to impute such a capability to our probes, we use the probes we learned above in a conformal prediction framework. Specifically, we use a held-out validation set $\mathbb{D}_{\text{valid}}$ to calibrate the probe after training. We compute the probes’ prediction probabilities $\mathbf{\hat{y}}\in[0,1]^{|c|}$ for each class with true label $y\in\{1,\ldots c\}$ on the validation set, and find the lowest threshold quantile $q$ that satisfies:
@@
During inference, when the probe predicts a behavior with probability vector $\mathbf{\hat{y}}_{\text{test}}$, we return the prediction if and only if $\max{(\mathbf{\hat{y}}_{\text{test}})}\geq q$, otherwise we defer to the LM. In all experiments we set $\alpha=0.9$. If no satisfying $q$ can be found we defer on all test instances. We ablate $\alpha$ in Section 6.
Using the conformal probes (Figure 2) significantly increases the performance across all datasets. For datasets where the total test consistency was already high (ARC with $88\%$ total test consistency), the conformal probes predict on large portions of the test set with consistency close to 90%. On datasets where total test consistency was poor (MedMCQA, MMLU, PiQA), the probes predict on fewer instances. When they do predict, they do so with higher consistencies than on the full test set.
4 Accelerating Inference-Scaling with Conformal Probes
Figure 3: Tradeoff between accuracy loss (CoT Accuracy - Accelerated CoT Accuracy) and inference cost when using our method to accelerate CoT Prompting in IID (Left) and OOD (Right) settings. In IID setting, Accuracy loss is negative on average, and never more than 1.34%, while inference cost reductions are 65% on average. In the OOD setting, inference cost reductions are higher at 81% on average, with a small accuracy loss of 2.3% on average.
@@ 4 Accelerating Inference-Scaling with Conformal Probes
We extend the method described above to 18 other datasets spanning the tasks of Sentiment Analysis, Topic Analysis, Toxicity Detection and Fact Verification. Details on dataset setup are in Appendix C.
@@
For all of these, we use Llama3.1-8B under CoT prompting (outputting an explanation before the final answer). We train a linear probe that uses the internal representation of the final input token at the 18th layer to predict the class that the CoT model will eventually output, and then perform conformal calibration. During inference, if the conformal probe is confident we use the probe estimation as the final answer. If not, we allow the model to continue its CoT generation and provide the final answer. We compare this to the vanilla CoT on two metrics—Accuracy Loss (Accuracy of CoT - Accuracy of Method) and Inference Cost Reduction (#CoT Forward Passes - #Method Forward Passes# CoT Forward $\frac{\#\text{CoT Forward Passes - \#Method Forward Passes}}{\text{\# CoT Forward Passes}}$).
The results (Figure 3) show that the method is highly effective at reducing the inference cost with minimal cost to accuracy. The minimum inference cost reduction is $4.7\%$, with an average reduction of $65\%$ across all datasets. Despite this significant speedup, the average accuracy loss is near zero ($\mathbf{-0.46\%}$), with the worst loss at $1.34\%$. Surprisingly, accuracy increases on several datasets.
@@
Finally, we investigate the out-of-distribution generalization capabilities of the conformal probes. For each test dataset of the MCQ and Sentiment tasks, we train and calibrate using data from every other dataset. The results (Figure 3) show that the probes do exhibit OOD generalization, suggesting the method may be applicable even when there are slight shifts between training and test distributions.
5 Creating Early Warning Systems for Behaviors of Interest
(a) Unanswerable Question Abstention
(b) Bullet Point Format Following
(c) JSON Format Following
(d) Perplexity Confidence Estimation
(e) Verbalized Confidence Estimation
Figure 4: Probes can provide precise early warning signals for a range of important behaviours
The phenomenon of probes containing information on the eventual output sequence of the LM is not limited to predicting text classification answers but applies to a wide range of behaviors that can be represented with discrete labels. In this section, we showcase several other use cases of this finding. We describe the key experiment designs below, with exact prompts in the Appendix C.
LM Abstention: We collect two datasets (SelfAware [68], KnownUnknown [4]) which have both answerable and unanswerable questions. The LM is guided, via few-shot examples, to first reason out loud with respect to the question content and then either answer it or refuse to answer if the question seems to be unanswerable. Separately, we collect malicious and benign prompts from WildJailbreak [30], and prompt the LM to refuse to comply with a request if it is malicious. In both cases, the conformal probes try to estimate whether the LM will give an answer or refuse.
Results show that probes confidently estimate (Figure 4(a)) the abstention decision on over 15% of KnownUnknown, 44% of SelfAware and 83% of WildJailbreak, with over 90% across datasets.
Format Following Errors: We collect 3 QA datasets (NaturalQA [38], MSMarco [46], TriviaQA [32]) and prompt the LM to an answer in a specific format. One format requires the answer to be organized in exactly 3 bullet points, and the other requires a JSON output which contains pre-specified fields. The probes estimate whether the output will adhere to the format.
Predicting format following errors in a JSON format proves more challenging than the bullet point format( 4(b)). However, even on the JSON task, the probes maintain a minimum consistency of 84% and have confidence on over 8% of instances on 2/3 datasets while deferring completely on the other.
Low Confidence Output: On the same 3 QA datasets, we try to identify when the LM is less confident in its response. Following recent work [33, 31], we consider one setting where we explicitly prompt the model to verbalize its confidence, and another where we use the per-token perplexity of the output as a proxy for the confidence. In the case of the perplexity measure, we consider the bottom $25\%$ of scores to be ‘high confidence’, and the top $25\%$ to be ‘low confidence’ (discarding the rest). In both cases, the probes attempt to identify whether the model will be confident in its output.
This is the most challenging task, probes typically make predictions on around 4% of test instances. However, consistency remains high, making the early warning system feasible in practice.
The conformal probe-based early warning systems vary in their confidence across tasks and datasets. However, on a wide range of LM behaviors, they predict on sizable portions of the test set with consistently high consistency.

### v2

@@ Abstract
The text produced by language models (LMs) can exhibit specific ‘behaviors,’ such as a failure to follow alignment training, that we hope to detect and react to during deployment. Identifying these behaviors can often only be done post facto, i.e., after the entire text of the output has been generated. We provide evidence that there are times when we can predict how an LM will behave early in computation, before even a single token is generated. We show that probes trained on the internal representation of input tokens alone can predict a wide range of eventual behaviors over the entire output sequence. Using methods from conformal prediction, we provide provable bounds on the estimation error of our probes, creating precise early warning systems for these behaviors. The conformal probes can identify instances that will trigger alignment failures (jailbreaking) and instruction-following failures, without requiring a single token to be generated. An early warning system built on the probes reduces jailbreaking by 91%. Our probes also show promise in pre-emptively estimating how confident the model will be in its response, a behavior that cannot be detected using the output text alone. Conformal probes can preemptively estimate the final prediction of an LM that uses Chain-of-Thought (CoT) prompting, hence accelerating inference. When applied to an LM that uses CoT to perform text classification, the probes drastically reduce inference costs (65% on average across 27 datasets), with negligible accuracy loss. Encouragingly, probes generalize to unseen datasets and perform better on larger models, suggesting applicability to the largest of models in real-world settings.  [^1] Our code is accessible at: https://github.com/DhananjayAshok/LMBehaviorEstimation
@@ 1 Introduction
Language models (LMs) have emerged as the dominant approach to general language tasks [88, 22], seeing widespread adoption in chatbots [3, 68], code generation [18] and reasoning systems [67]. However, they have been known to ‘misbehave,’ hallucinating false information [98], failing to adhere to output specifications [59], or, most concerningly, producing content that is misaligned with human values [75], sometimes dangerously so [35]. In practice, such misbehavior is mitigated by deploying post-hoc guardrails on the output of the LM [92, 24, 16, 101, 8, 53]. For example, a chatbot LM may exhibit harmful behavior by complying with a malicious request [41, 90]; developers will typically attempt to detect this behavior in the output [101] and return an abstention message instead of the original LM content. Such a framework suffers the economic [100] and environmental [83, 76] expense of generating every output token, a cost that threatens to grow exponentially as frontier labs continue to scale both model sizes [37, 49] and the number of tokens generated during inference [91, 95, 67, 32].
@@
In this paper, we show that the internal representations of LMs can be used to preemptively predict behaviors that will emerge over the entire output sequence. This information becomes accessible before the LM has generated a single token, enabling the system to take appropriate actions without suffering inference costs on those instances. Concretely (Figure 1), we train linear classifiers (probes) [4] that use an LM’s internal representation of input tokens to predict the eventual behavior of its output. We then calibrate the probes using methods from conformal prediction [81]. During inference, we look to probes to make a prediction only when there is a provable guarantee on the estimation error, ensuring precise early warning signals for various model behaviors.
We show that these probes can preemptively identify degenerate behavior, such as instruction-following failures (Section 3). Across multiple datasets and output formats, our probes give precise early warning signals for cases where the LM will fail to follow specified output format instructions.
Conformal probes can also detect deeper, ‘self-reflective’ properties, like whether an LM’s behavior conflicts with its safety alignment. We demonstrate this (Section 4.1) by applying our method to enhance the safety of an LM that is instructed to abstain from complying with malicious requests. Probes are trained to identify cases where this will not happen—the request will be malicious and the model will not abstain. Once again, our probes are able to precisely detect (with over 92% accuracy) cases where the LM complies with malicious requests. An early warning system built on these conformal probes reduces the jailbreak success rate from 30% to 2.7%, a 91% reduction.
We further show that conformal probes can identify behaviors that cannot be measured with output text. We show (Section  4.2) that our probes can a priori estimate an LM’s confidence as measured by the per-token perplexity of the output; as this model-specific property cannot be inferred from just the text sequence alone, we show the wide scope of applicability of our probing approach.
@@
Finally, we apply our method to an LM that tackles text classification by generating several explanation tokens before a class prediction, i.e. using Chain-of-Thought (CoT) reasoning [91] (Section 5). The probes estimate the final class prediction that will appear after the reasoning chain and exit early if confident in their estimate. On 27 datasets, spanning the tasks of Multiple Choice QA, Sentiment Analysis, Topic Classification, Toxicity Detection and Fact Verification, our method reduces inference costs by 65% on average, while suffering accuracy losses of no more than 1.4% (worst case). The probes generalize to unseen datasets, reducing the inference cost of CoT on OpenBookQA [63] by 68% with minimal loss to accuracy (0.4%), despite only training on other MCQA datasets.
@@
We demonstrate that often fewer than 500 training instances are sufficient for the probes to attain high estimation consistency, while ablations on the probing layer present a more complicated, task-specific story. Encouragingly, increasing the scale of the LM improves the performance of our technique, suggesting it may scale favorably with ever-increasing model sizes [15, 37, 19, 26].
@@
Figure 1: Overview of our method as applied in preemptive safety alignment failure detection. During training, we use many input prompts to collect a dataset consisting of the hidden states of the input tokens and eventual output behavior (in this example, the behavior is whether the LM fails to abstain on a malicious prompt). The dataset is used to train a probe on the hidden states. With a validation set, we calibrate the probes for conformal prediction by identifying a threshold for confident predictions. During inference, if the probe is confident in its estimate of the eventual LM behavior, we stop computation early (in this example, overriding the LM and abstaining).
@@ 2 Background
Hidden State Probing: Lightweight probes have long been used to interpret the internal activations of neural networks [4] and language models [71, 10]. Given a set of input prompts, we compute the generated outputs and store the internal activations computed during the forward passes of an LM (e.g. the outputs of a specific Transformer [88] layer). The instances are then assigned a classification label meant to capture some property of the input and output. The probes are then trained to predict the label using only the internal activations as input. For example, Azaria and Mitchell [10] collect the activations of the middle layer of an LM when processing the output tokens, and manually label the generations as true or false. They then train a small neural network which can identify whether the model output is truthful, given the internal activations of the generated tokens. Prior work in probing uses the internal states of all tokens, even generated ones, which incurs the cost of inference. In this work, we explore an alternate approach, training probes on the internal states of the input tokens alone, allowing us to predict properties of the output tokens before they have been generated.
@@
At inference time, the prediction set for a test instance $\mathbf{x}$ is $\{\hat{c}|S(\mathbf{x},\hat{c})\leq q_{\alpha}\}$. For the above score function, no two classes can score greater than 0.5 simultaneously, implying that with $\alpha>0.5$ the prediction set at inference time is either empty (deferral/abstention) or consists of a single class label.
@@ 3 Can Internal States Predict Eventual Behavior?
In this section, we ask whether the internal representations of the LMs input can be used to predict eventual behavior, even before any output token has been generated. We explain our methodology with the running example of an LM instructed to answer a question in bullet point format.
The LM is expected to provide an answer in two bullet points, with any fewer or greater bullets in the answer being a failure to follow instructions. Suppose we are given an input prompt:
   Question: What do Osteoclasts do?
A LM may respond to this with:
1. Decompose bone 2. Differentiate from monocyte 3. Create space for tissue
We can identify that the LM has behaved contrary to its instructions when it generates the ‘3’. The previous tokens give no indication that a failure was imminent; however, is it possible that the model contains signals that can identify this eventual failure by the time it generates the first token, i.e. ‘1’?
[…]
To test this, we collect three QA datasets—NaturalQA [54], MSMarco [65] and TriviaQA [47] and prompt the LM to answer questions in a specific format. One format requires the answer to be organized in exactly three bullet points, while the other requires a JSON output with pre-specified fields. We evaluate whether the output format has been followed, and then sample this data to obtain training and testing splits that have an equal number of failures and successes.
We collect the output of the middle Transformer layer of Llama3.1-8B [26] when processing the final token of the input (in the example above, the ‘?’), then train linear classifiers to predict whether the output will eventually fail to follow instructions. We ablate layers in Section 6, reproduce results on other LMs in Appendix A, and provide the prompts that specify the format in Appendix C.
We also fine-tune a bert-large-uncased [22] model for text classification, which involves tuning 340M parameters, $4\times 10^{4}$ times the number of trainable parameters of the linear model. The BERT model is trained to predict whether the output will fail to follow formatting instructions given the text of the input prompt alone, and is a measure of how a significantly stronger model performs when attempting to establish a pattern between the inputs to the model and the output behavior. Methods are evaluated on the metric of estimation consistency, i.e, how accurate they are at preemptively predicting whether or not the formatting instructions will be followed.
The probes (Figure 2, navy blue bar) outperform the random baseline in all settings, showing that an LM’s representation of the input tokens contains information on behavior that will emerge over the entire output sequence. Moreover, the probes also outperform fine-tuning on BERT in all but one of the settings, despite using around only 0.0025% as many parameters.
However, despite outperforming the baselines, the average estimation consistency of the probes is less than 75%. A naive attempt to act on the probes estimate would result in catastrophic failure.
@@ 3.1 Creating robust behavior estimators with conformal prediction
Internal states may occasionally be insufficient to estimate eventual behavior. The ideal system handles such cases, making consistent predictions when confident and deferring otherwise. Hoping to impart such a capability to our probes, we use the probes learned above in a conformal prediction framework. Specifically, we use a held-out validation set $\mathbb{D}_{\text{valid}}$ to calibrate the probe after training. We compute the probes’ prediction probabilities $\mathbf{\hat{y}}\in[0,1]^{|c|}$ for each class with true label $y\in\{1,\ldots c\}$ on the validation set, and find the lowest threshold quantile $q$ that satisfies:
@@
During inference, when a probe predicts a behavior with probability vector $\mathbf{\hat{y}}_{\text{test}}$, we return the prediction if and only if $\max{(\mathbf{\hat{y}}_{\text{test}})}\geq q$, otherwise we defer to the LM. Unless specified otherwise, we set $\alpha=0.9$. If no satisfying $q$ can be found, we defer on all instances. We ablate $\alpha$ in Section 6.
Using conformal probes (Figure 2) significantly increases performance across all datasets. At $\alpha=0.8$, the probes have a test coverage of 40% and, for the covered datapoints, estimation consistency is 80.7%, showing that the conformal probes are well calibrated to the user-specified confidence level. Raising $\alpha$ to $0.9$ results in a reduced average coverage of 10%, but an increased average consistency of 89.2%. This allows users to flexibly trade off between coverage and consistency, enabling precise and trustworthy early warning systems that do not offer an opinion when confidence is low.
Conformal Probe ($\alpha=0.9$) BERT
Dataset Coverage Consistency Prec. Rec. Consistency Prec. Rec.
SelfAware 100.0 98.0 94.5 94.5 93.3 91.7 84.4
KnownUnkown 92.3 92.6 75.1 72.8 67.1 64.4 93.1
WildJailbreak 100.0 94.0 90.0 91.0 83.4 73.8 33.3
Table 1: Probes estimate whether the LM will fail to abstain when it should have with significantly higher precision and recall than a BERT baseline. On WildJailbreak, where 30% of malicious prompts are mistakenly complied with, using the probe as an early warning signal reduces successful jailbreaks to 2.7%. In comparison, using the BERT model would leave jailbreak success rate at 20%.
4 Creating Early Warning Systems for Other Behaviors of Interest
In this section, we demonstrate that probes are not limited to preemptively identifying degenerate behavior, such as failing to follow instructions. We show that conformal probes can serve as early warning systems for instances that will trigger failures to follow safety alignment, and those that will result in the model giving a low-confidence response. We describe the key experiment designs below, with exact prompts and details in the Appendix C.
4.1 Detecting failures to follow safety alignment
The deployment of LM-powered chatbots raises concerns of harmful behavior that is unaligned with human interests [28]. One common setting where such behaviors arise is when the user requests information that does not exist, with models often hallucinating an answer regardless, misleading the user in the process [9]. Another common occurrence is when the user themselves requests the model to help them complete malicious tasks, such as building a weapon of mass destruction [90].
We collect two datasets (SelfAware [96], KnownUnknown [5]), which have answerable and unanswerable questions. The LM reasons as to whether the question is answerable, providing a response if it is and abstaining otherwise. Similarly, we collect malicious and benign prompts from WildJailbreak [45], and prompt the LM to refuse to comply with a request if it is malicious. Inspired by work in AI safety and alignment, we adopt the viewpoint that answering an unanswerable question or complying with a malicious request is disproportionately problematic [11]. We filter our training set to only include the instances where the LM answers or complies with the request, and train probes to identify the cases where the LM should have abstained (i.e. fails to avoid answering an unanswerable question or falls victim to jailbreaking). During inference, if the probe is confident ($\alpha=0.9$) that the LM has failed to abstain when it should have, we override the LM and abstain instead.
Results show (Table 1) that probes greatly enhance the safety of the LMs, outperforming the BERT baseline on consistency, precision and recall. Across all datasets, probes identify over 72% of the failures to abstain (over 90% for 2/3 datasets). This safety comes at little cost; when the probes force an abstention, they are correct over 86% of the time (on average). They make confident estimations on at least 92% of the samples, and maintain conformal estimation consistency of at least 92%. Concretely, on WildJailbreak the use of conformal probes reduces the percentage of successful jailbreaks from 30% to 2.7%, a significant safety improvement.
4.2 Quantifying LM uncertainty
Figure 3: Estimation consistency when preemptively predicting an LM’s confidence in its response. Numbers inside the bars in bold italics are the coverage for conformal probes that engage in selective prediction. Probes outperform all other methods, with the confidence estimation tasks proving more challenging for BERT. Conformal probes have lower coverage rates than the format following (Section 3) and safety alignment (Section 4.1) tasks. However, they maintain consistency at the user-defined confidence level, ensuring that their early warning signals are reliable.
Language models are prone to hallucination [98] and often have gaps in their knowledge [38], making methods that estimate the trustworthiness of specific outputs vital. Recent work [48, 46] approaches this problem by using the model itself to quantify how certain it is in its answer, either by explicitly prompting it to verbalize its confidence or using the per-token perplexity of the output. If we detect an untrustworthy response, we may abstain from answering or escalate the query to a more capable model. An early warning system that preemptively detects such responses before they are even generated would allow us to take these actions more efficiently.
We train conformal probes to preemptively estimate how confident the LM will be in its response using both of these uncertainty quantification methods on NaturalQA, MSMarco and TriviaQA. In the case of the perplexity measure, we consider the bottom $25\%$ of scores to be ‘high confidence,’ and the top $25\%$ to be ‘low confidence’ (discarding the rest). In both cases, the probes attempt to identify whether the model will be confident in its output. This task proves more challenging for all methods, however, probes continue to outperform both the random baseline and the much larger BERT model. Despite this, the conformal probes maintain high consistency, making the early warning and redirection system feasible in practice.
5 Selective Inference-Scaling with Conformal Probes
@@
In this section, we show how our method can be used to accelerate systems that use CoT prompting with an LM to perform text classification. We collect 27 text classification datasets, spanning the tasks of Multiple Choice Question Answering, Sentiment Analysis, Topic Analysis, Toxicity Detection and Fact Verification. Details on dataset setup are in Appendix C.
@@
We use Llama3.1-8B under CoT prompting to perform text classification, i.e., outputting an explanation before the final class prediction. We train a linear probe that uses the internal representation of the final input token at the 18th layer (based on results from previous experiments) to predict the class that the CoT model will eventually output, and then perform conformal calibration. During inference, if the conformal probe is confident, we interrupt generation and use the probe estimation as the final answer. If not, we allow the model to continue its CoT generation and provide the final answer. We note that this method of acceleration is synergistic with other inference optimization algorithms such as Speculative Decoding [55] or architectural changes [27]. This is because in the cases where the probe does not interrupt the LM, it does not influence the remaining forward passes, hence conformal probes can be flexibly added on top of any inference paradigm for further efficiency.
We compare to vanilla CoT on two metrics—Accuracy Loss (Accuracy of CoT - Accuracy of Method) and Inference Cost Reduction (#CoT Forward Passes - $\frac{\#\text{CoT Forward Passes - \#Method Forward Passes}}$$\text${\# CoT Forward Passes}}).
Figure 4: Tradeoff between accuracy loss and inference cost when using probes to accelerate CoT Prompting in IID and OOD settings. Using IID probes leads to a 1.34% accuracy loss (worst case) and a negative accuracy loss on average (i.e. accuracy improves). Inference cost reductions are 65% on average. Probes generalize to OOD data, inference cost reductions are higher at 81% on average, with a small accuracy loss of 2.3% on average.
@@
The results (Figure 4, in blue) show that the method is highly effective at reducing the inference cost with minimal cost to accuracy. The minimum inference cost reduction is $4.7\%$, with an average reduction of $65\%$ across all datasets. Despite this significant speedup, the average accuracy loss is near zero ($\mathbf{-0.46\%}$), with the worst loss at $1.34\%$. Surprisingly, accuracy increases on several datasets. Finally, we investigate the out-of-distribution generalization capabilities of the conformal probes. For each test dataset of the MCQ and Sentiment tasks, we train and calibrate using data from every other dataset. The results (Figure 4, in orange) show that the probes do exhibit OOD generalization, suggesting the method may be applicable even when there are slight shifts between training and test distributions.